OpenStudy (anonymous):

What is the domain of x in this expression? y+2/x+3 + √16x A. all real numbers B. all real numbers except -3 and -2 c. all real positive numbers and 0 D. all real numbers except -3,-2, and 0

OpenStudy (anonymous):

Can you rewrite in the equation editor? Your syntax is a little unclear, but is it: \[\frac{y+2}{x+3}+\sqrt{16x}\]

OpenStudy (anonymous):

yes exactly like that

OpenStudy (anonymous):

is it y+2 or y=?.... should be y(x)=

OpenStudy (anonymous):

y+2

OpenStudy (anonymous):

Okay, domain is defined as possible x-values, right? so what x value would give us an undefined value or an imaginary number?

OpenStudy (anonymous):

0?

OpenStudy (anonymous):

not quite. if x+3=0, then our first term (y+2)/(x+3) becomes undefined. Also if x is a negative number in our second term sqrt(16x) then we get an error because you can't take the square root of a negative number, right?